How do I completely factor the expression:
w^2 - 15 w +54

Thank you for your help.

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Find a pair of factors of 54 with a sum of 15.
---------
PS Both are negative.

Answer by ikleyn(52894)

.
How do I completely factor the expression:
w^2 - 15 w +54
~~~~~~~~~~~~~~~~~~~~~

54 = 6*9 = (-6)*(-9),  and

-15 = (-6) + (-9).


Therefore,

w^2 - 15 w +54 = (x-6)*(x-9).